Diet Methods: Here we study the effectiveness of two dieting methods. Method 1 is a diet high in protein, fiber, and fat but low in carbohydrates. Method 2 is a diet low in fat but high in fiber and carbohydrates. After one year, a sample of participants from each diet method were checked for the amount of weight lost. The 95% confidence intervals for the mean weight loss of each plan are given below.

Method 1:   8.87 < μ < 9.73
Method 2:   7.16 < μ < 8.44

(a) What is the sample mean weight loss for participants in the Method 1 diet plan? Round your answer to 1 decimal place.

x1 =  pounds



(b) What is the sample mean weight loss for participants in the Method 2 diet plan? Round your answer to 1 decimal place.

x2 =  pounds



(c) Are you 95% confident that the mean weight loss for all those on the Method 1 diet plan is greater than the mean weight loss for all on the Method 2 diet plan and why?

No, because the confidence intervals overlap.Yes, because upper limit for Method 2 is less than the lower limit for Method 1.     No, because upper limit for Method 2 is less than the lower limit for Method 1.Yes, because the confidence intervals overlap.

Additional Materials
Answer #1

a)
mean, x1bar = (9.73 + 8.87)/2 = 9.3 pounds

b)
mean, x2bar = (8.44 + 7.16)/2 = 7.8 pounds

c)
Yes, because upper limit for Method 2 is less than the lower limit for Method 1.